# Combining Crown Structures for Vulnerability Measures

### v1

@@ 3.1 Vertex Integrity
Note that the lower bound of $\lambda+1$ in Lemma 3.5 is applicable to the induced graph of $V^{\prime}=C\cup H\cup V(\mathfrak{R})$ within $G$. This implies that the isolated components of $G-V^{\prime}$, defined as having a size (or weight) of at most $\lambda$, can be safely removed. Their removal does not affect the decision problem of VI or wVI concerning $p$.
Clearly, a yes-instance cannot have a lower bound $p_{\ell}$ larger than $p$ as stated in the following corollary.
@@
Let $(G,p)$ be an instance of the vertex integrity problem and for $\lambda\in[p]$ let $(C,H,\mathfrak{R},f)$ be a $\lambda$-BCD $(C,H,\mathfrak{R},f)$ in $G$ with $|H|+|\mathfrak{R}|\leq p$. If $|C|+|H|+|V(\mathfrak{R})|\geq 3p^{2}$, then $G^{\prime}=G[C\cup H]$ contains a $(p,\lambda)$-CD in $G$. Furthermore, we can extract it from $G^{\prime}$ in time $\mathcal{O}(|V(G^{\prime})||E(G^{\prime})|)\subseteq\mathcal{O}(|V(G)||E(G)|)$.
If we now combine Lemma 3.5 and Lemma 3.7, we can finally use Lemma 3.4 by finding a $\lambda\in[p-1]$ such that a ($\lambda+1$)-BCD $(C,H,\mathfrak{R},f)$ satisfies $|H|+|\mathfrak{R}|\leq p$, while a $\lambda$-BCD $(C^{\prime},H^{\prime},\mathfrak{R}^{\prime},f^{\prime})$ satisfies $|H^{\prime}|+|\mathfrak{R}^{\prime}|>p$, which only adds a factor of $\mathcal{O}(\log(p))$ to the computation cost. Formulated more precisely: For $Q\in\mathbb{CC}(C)$ we have $|Q|\leq\lambda+1\leq p_{\ell}$ as $\lambda<p_{\ell}$ by Lemma 3.5. Further, we remove all vertices that are in components of size at most $\lambda$, which is a save reduction rule as explained above. If the vertex size of the remaining input graph, i.e. $G[C\cup H\cup V(\mathfrak{R})]$ is at least $3p^{2}$, then we can find a ($p,p_{\ell}$)-CD for $G$ in $G[C\cup H]$ in polynomial time by Lemma 3.7, which is a reducible structure by Lemma 3.4.
With these facts in hand we can design a kernelization algorithm. It takes as input an instance $(G,p)$ of VI and returns an equivalent instance of with at most $3p^{2}$ vertices.
Find reducible structures (AlgVI):
@@
    Compute a $p$-BCD $(C_{1},H_{1},\mathfrak{R}_{1},f_{1})$. If $|H_{1}|+|\mathfrak{R}_{1}|>p$ return a trivial no-instance.
@@
    Compute a $1$-BCD $(C_{2},H_{2},\mathfrak{R}_{2},f_{2})$. Let $V^{\prime}=V(G)\setminus(C_{2}\cup H_{2}\cup V(\mathfrak{R}_{2}))$. If $|H_{2}|+|\mathfrak{R}_{2}|\leq p$ and $V^{\prime}\neq\varnothing$, then return $(G-V^{\prime},p)$. If $|H_{2}|+|\mathfrak{R}_{2}|\leq p$ and $V^{\prime}=\varnothing$, then compute a $(p,1)$-CD ($H_{2}^{\prime}$,$C_{2}^{\prime}$,$f_{2}^{\prime}$) in $G[H_{2}\cup C_{2}]$ and return $(G-\{H_{2}^{\prime}\cup C_{2}^{\prime}\},p-|H_{2}^{\prime}|)$.
@@
    Otherwise, for $\lambda\in[p-1]$ find a ($\lambda+1$)-BCD $(C,H,\mathfrak{R},f)$ and a $\lambda$-BCD $(C^{\prime},H^{\prime},\mathfrak{R}^{\prime},f^{\prime})$, such that $|H|+|\mathfrak{R}|\leq p$ and $|H^{\prime}|+|\mathfrak{R}^{\prime}|>p$. Let $V^{\prime\prime}=V(G)\setminus(C\cup H\cup V(\mathfrak{R}))$. If $V^{\prime\prime}\neq\varnothing$, then return $(G-V^{\prime\prime},p)$. Otherwise, compute a $(p,\lambda+1)$-CD ($H^{\prime\prime}$,$C^{\prime\prime}$,$f^{\prime\prime}$) in $G[C\cup H]$ for $G$ and return $(G-\{C^{\prime\prime}\cup H^{\prime\prime}\},p-|H^{\prime\prime}|)$.
We already explained why we can safely remove $V^{\prime}$ or $V^{\prime\prime}$ in step 2 and 3 if these are not empty. Note that if $V^{\prime}$ or $V^{\prime\prime}$ is empty, then the vertex sets of the corresponding BCD’s, i.e. $C_{2},H_{2},V(\mathfrak{R}_{2})$ and $C,H,V(\mathfrak{R})$, respectively, form a partition of $V(G)$. The correctness of step 1 is implied by Corollary 3.6. Step 2 is correct because $p_{\ell}\geq 1$ and if $V^{\prime}=\varnothing$ we obtain by Lemma 3.7 and $|V|=|C_{2}|+|H_{2}|+|V(\mathfrak{R}_{2})|\geq 3p^{2}$ that there must be a $(p,1)$-CD if $|H_{2}|+|\mathfrak{R}_{2}|\leq p$. Analogously, for step 3 as $p_{\ell}\geq\lambda+1$ by Lemma 3.5 in this step. For the overall correctness of algorithm AlgVI, it remains to be proven that if we reach step 3, the required $\lambda$ exists. For a $\lambda$-BCD in a binary search, we cannot ensure that $|H|+|\mathfrak{R}|$ increases with decreasing $\lambda$ values, because the sizes of $H$ and $\mathfrak{R}$ are not necessarily monotonic with respect to $\lambda$. Note, for example, that the elements in $\mathfrak{R}$ have a size range from $\lambda+1$ to $3\lambda$. For a successful binary search, however, it is sufficient to know that there is a $\lambda$ that satisfies the desired properties for step 3. Since we only enter this step if the extreme cases ($p$-BCD and 1-BCD) in step 1 and 2 hold, there has to exist at least one such $\lambda$ value in-between.
Finally, we prove the specified running time of Theorem 3.1, which completes the proof of Theorem 3.1. Note that for a yes-instance we have to call the algorithm AlgVI at most $|V|$ times to guarantee the desired kernel.
@@ 3.2 Weighted Vertex Integrity and Component Order Connectivity
In this section, we shift our focus to the weighted variants, namely wVI and wCOC. While utilizing the packing size of the associated BCD offers a starting point for deriving a lower bound for VI, it proves insufficient for improvements in the weighted setting. Therefore, we integrate the weight of the separator $H$ within a BCD $(C,H,\mathfrak{R},f)$ into our analysis. Additionally, we incorporate two distinct DBE’s into a BCD in case a reduction is not achievable within the respective setting. This approach enables us to establish a tighter lower bound for $p_{\ell}$, estimate the remaining instance size more accurately to obtain the desired kernelization results (cf. Theorems 3.2 and 3.3), or a identify no-instance.
For improved readability, we extend the notation of the inverse function $f^{-1}$. Henceforth, sometimes it returns a vertex set that is the union of the corresponding vertex sets, which is always clear from the context. For instance, for $h\in H$ or $H^{\prime}\subseteq H$ we have $f^{-1}(h)=V(f^{-1}(h))$ and $f^{-1}(H^{\prime})=V(f^{-1}(H^{\prime}))$, respectively. Let us start by explaining the type of reducible structure we are searching for.
@@
Let $G=(V,E,w)$ be a vertex weighted graph, $a,b\in\mathds{N}$, $(H,H^{\prime},b,f,D)$ a DBE in $G$ with $D=\{d_{h}\}_{h\in H}$, where $d_{h}=a-2+b\cdot(w(h)+1)$ for each $h\in H$, and let $C=f^{-1}(H)$.
  1. 1.
    Let $(G,p)$ be an instance of the weighted vertex integrity problem, $a=p$ and $1\leq b\leq p_{\ell}$. Then, $(G,p)$ is a yes-instance if and only if $(G-\{H\cup C\},p-w(H))$ is a yes-instance.
  2. 2.
    Let $(G,k,W)$ be an instance of the weighted component order connectivity problem and $a,b=W$. Then, $(G,k,W)$ is a yes-instance if and only if $(G-\{H\cup C\},k-w(H),W)$ is a yes-instance.
We are now introducing the two DBE’s mentioned in conjunction with a BCD. Let $G=(V,E,w)$ be a vertex weighted graph, $a,s,\lambda\in\mathds{N}$, $(C,H,\mathfrak{R},f)$ a $\lambda$-BCD with $\lambda\in[a]$ and $|H|+|\mathfrak{R}|\leq s$, where $\max_{v\in V}w(v)\leq\max(a,s)=:\mu$. Let $D^{Y}=\{d^{Y}_{h}\}_{h\in H}$ and $D^{Z}=\{d^{Z}_{h}\}_{h\in H}$ be demands, where $d^{Y}_{h}=a-2+\lambda\cdot(w(h)+1)$ and $d^{Z}_{h}=w(h)-1+(\sqrt{s}+1)\lambda$ for each $h\in H$. Let $(Y_{1},Y_{2},\lambda,f_{Y},D^{Y})$ and $(Z_{1},Z_{2},\lambda,f_{Z},D_{Z})$ be DBE’s in $G[C\cup H]$ with $Y_{1},Y_{2},Z_{1},Z_{2}\subseteq H$ and $f_{Y}^{-1}(Y_{1}),f_{Y}^{-1}(Y_{1}),f_{Z}^{-1}(Z_{1}),f_{Z}^{-1}(Z_{2})\subseteq\mathbb{CC}(C)$. Observe that if $Y_{1}\neq\varnothing$ (resp. $Z_{1}\neq\varnothing$) then this set separates $f^{-1}_{Y}(Y_{1})$ (res. $f^{-1}_{Z}(Z_{1})$) from the rest of the graph, since $H$ separates $C$ from the rest of the graph. Thus, if $a=p$ and $\lambda\leq p_{\ell}$ considering wVI or $a,\lambda=W$ considering wCOC and $Y_{1}\neq\varnothing$, then in both cases we would have a reducible structure for the corresponding problem (cf. Lemma 3.9). We conclude this section by presenting two crucial lemmas (the continuation can be found in LABEL:appendix::continuation-wVI-wCOC). These lemmas serve as the foundation for designing the algorithms needed to prove Theorems 3.2 and 3.3. The first lemma aims to estimate the instance size, while the second lemma helps establish a more precise lower bound for $p_{\ell}$ or identify instances with no feasible solutions.
Lemma 3.10.
Let $w(Z_{1})\leq 2\mu^{1.5}$. If $Y_{1}=\varnothing$, then $w(V)<3\mu(s+\sqrt{\mu}\lambda)$.

### v2

@@ 3.1 Vertex Integrity
Proof 3.11.
Let $V^{\prime}=C\cup H\cup\mathfrak{R}^{\prime}$ and consider the graph $G[V^{\prime}]$ (subgraph of $G$ created by removing all components of weight at most $\lambda$). The vertex sets $\mathfrak{R}\cup\{\{h\}\cup V(f^{-1}(h))\}_{h\in H}$ form a $(\lambda+1)$-packing of size $|\mathfrak{R}|+|H|>p$ in $G[V^{\prime}]$. It follows that we need more than $p$ vertices for a vertex separator $S$ in $G[V^{\prime}]$, such that $\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq\lambda$. Thus for any $U\subset V$ with $|U|\leq p$ it follows that $\max_{Q\in\mathbb{CC}(G-U)}w(Q)>\lambda$. This implies $\lambda<p_{\ell}$ for the instance $(G[V^{\prime}],p)$, since any solution has cardinality at most $p$.
An additional advantage of a $(\lambda+1)$-BCD $(C,H,\mathfrak{R},f)$ is that the balanced part $\mathfrak{R}$ can be upper bounded by $3(\lambda+1)|\mathfrak{R}|\leq 3p_{\ell}|\mathfrak{R}|<3p|\mathfrak{R}|$, while a suitable $\lambda$ choice also upper bounds $|\mathfrak{R}|$ by $p$. In particular, if $H=C=\varnothing$ and $|\mathfrak{R}|\leq p$, then we would already have an instance with a size (or weight) of at most $3p^{2}$.
Clearly, a yes-instance cannot have a lower bound for $p_{\ell}$ larger than $p$ as stated in the following corollary.
@@
Let $(G,p)$ be an instance of VI and for $\lambda\in[p]$ let $(C,H,\mathfrak{R},f)$ be a $\lambda$-BCD in $G$ with $|H|+|\mathfrak{R}|\leq p$. If $|C|+|H|+|V(\mathfrak{R})|\geq p\max\{3\lambda,p+\lambda\}$, then $G^{\prime}=G[C\cup H]$ contains a $(p,\lambda)$-CD in $G$. Furthermore, we can extract it from $G^{\prime}$ in time $\mathcal{O}(|V(G^{\prime})||E(G^{\prime})|)\subseteq\mathcal{O}(|V(G)||E(G)|)$.
Proof 3.14.
Consider Theorem 1 on $G^{\prime}=G[H\cup C]$ (so $H$ in the role of $A$ and $C$ in the role of $B$), and with $d_{h}=p-1+\lambda$ and $y=\lambda$. This theorem then gives us a DBE $(H_{1},H_{2},\lambda,f,p-1+\lambda)$, with the guarantee $H_{1}\neq\emptyset$ in particular if $|H|+|C|\geq\sum_{h\in H}d_{h}$ (case $A=A^{\prime}$ in the Theorem). The part $(C_{1},H_{1},f)$ with $C_{1}=f^{-1}(H_{1})$ then gives us the desired non-empty $(H_{1},\emptyset,\lambda,f,p-1+\lambda)$ DBE, in other words a $(p,\lambda)$-CD in $G^{\prime}$. Moreover, observe that then $N(C)\subseteq H$ in $G$ and $H_{1}\subseteq H$, with $N(C_{1})\subseteq H_{1}$ in $G^{\prime}$ implies $N(C_{1})\subseteq H_{1}$ in $G$ and hence, $(C_{1},H_{1},f)$ is also a $(p,\lambda)$-CD in $G$.
Thus it remains to show $|H|+|C|\geq\sum_{h\in H}d_{h}=(p-1+\lambda)|H|$. Suppose $|C|+|H|<(p-1+\lambda)|H|$, then:
$|R|+|C|+|H|$ $<3\lambda|\mathfrak{R}|+(p-1+\lambda)|H|$
$\leq 3\lambda|\mathfrak{R}|+(p+\lambda)|H|$
$\leq\max\{3\lambda,p+\lambda\}(|\mathfrak{R}|+|H|)$
$\leq\max\{3\lambda,p+\lambda\}p,$
which is a contradiction to $|R|+|C|+|H|\geq\max\{3\lambda,p+\lambda\}p$.
It follows that $|C|\geq(p-2+\lambda)|H|$ and thus $|H|+|C|\geq|H|+(p-2+\lambda)|H|=\sum_{h\in H}d_{h}$. We can thus use Theorem 1 to extract a $(p,\lambda)$-CD $(C_{1},H_{1},f)$ with $\emptyset\neq H_{1}\subseteq H$ and $C_{1}\subseteq C$ from $G^{\prime}$ in time $\mathcal{O}(|V(G^{\prime})||E(G^{\prime})|)$.
If we now combine Lemma 3.10 and Lemma 3.13, we can finally use Lemma 6 by finding a $\lambda\in[p-1]$ such that a ($\lambda+1$)-BCD $(C,H,\mathfrak{R},f)$ satisfies $|H|+|\mathfrak{R}|\leq p$, while a $\lambda$-BCD $(C^{\prime},H^{\prime},\mathfrak{R}^{\prime},f^{\prime})$ satisfies $|H^{\prime}|+|\mathfrak{R}^{\prime}|>p$, which only adds a factor of $\mathcal{O}(\log(p))$ to the computation cost. Formulated more precisely: For $Q\in\mathbb{CC}(C)$ we have $|Q|\leq\lambda+1\leq p_{\ell}$ as $\lambda<p_{\ell}$ by Lemma 3.10. If the size of $G[C\cup H\cup V(\mathfrak{R})]$ is large, then we can find a ($p,\lambda$)-CD for $G$ in $G[C\cup H]$ in polynomial time by Lemma 3.13, which is a reducible structure by Lemma 6.
With these facts in hand we can design a kernelization algorithm. It takes as input an instance $(G,p)$ of VI and returns an equivalent instance with at most $3p^{2}$ vertices.
Find reducible structures (AlgVI)
@@
    Compute a $p$-BCD $(C_{0},H_{0},\mathfrak{R}_{0},f_{0})$. If $|H_{0}|+|\mathfrak{R}_{0}|>p$, return a trivial no-instance.
@@
    Compute a $1$-BCD $(C_{1},H_{1},\mathfrak{R}_{1},f_{1})$. If $|H_{1}|+|\mathfrak{R}_{1}|\leq p$, set $c=1$
@@
    Otherwise, for $\lambda\in[p-1]$ use binary search to find a ($\lambda+1$)-BCD $(C_{1},H_{1},\mathfrak{R}_{1},f_{1})$ and a $\lambda$-BCD $(C_{2},H_{2},\mathfrak{R}_{2},f_{2})$, such that $|H_{1}|+|\mathfrak{R}_{1}|\leq p$ and $|H_{2}|+|\mathfrak{R}_{2}|>p$. Set $c=\lambda+1$.
  4. 4.
    If $|C_{1}|+|H_{2}|+|V(\mathfrak{R}_{2})|\geq p\max\{3c,p+c\}$, compute a $(p,c)$-CD $(H^{\prime},C^{\prime},f^{\prime})$ in $G[C_{1}\cup H_{1}]$ and return $(G-(H^{\prime}\cup C^{\prime})+K_{c},p-|H^{\prime}|)$.
  5. 5.
    Delete all components of size at most $c$ and add a copy of $K_{c}$ in $G$.
Lemma 3.15.
For any instance $(G,p)$ with $p\geq 2$ and where not all components of $G$ have size at most $p$, exhaustive application of AlgVI produces an equivalent instance with at most $3p^{2}$ vertices.
Proof 3.16.
We first prove that an application of AlgVI is safe. The correctness of step 1 is implied by Corollary 3.12. Assuming $c$ satisfies $c\leq p_{\ell}$, the correctness of steps 5 and 4 follow directly from exhaustive application of Lemma 3.8 and Lemma 6, respectively.
For correctness, it remains to prove that $c$ satisfies $c\leq p_{\ell}$. Since the input $G$ is a graph with at least one component of size more than $p$, the choice $c=1$ in step 2 certainly satisfies $p_{\ell}\geq c$. For step 3 we first have to show that the claimed value for $\lambda$ with a ($\lambda+1$)-BCD $(C_{1},H_{1},\mathfrak{R}_{1},f_{1})$ and a $\lambda$-BCD $(C_{2},H_{2},\mathfrak{R}_{2},f_{2})$, such that $|H_{1}|+|\mathfrak{R}_{1}|\leq p$ and $|H_{2}|+|\mathfrak{R}_{2}|>p$ exists. For a $\lambda$-BCD computation in a binary search for the correct value for $\lambda$, we cannot ensure that $|H|+|\mathfrak{R}|$ increases with decreasing $\lambda$ values, because the sizes of $H$ and $\mathfrak{R}$ are not necessarily monotonic with respect to $\lambda$. (Note, for example, that the elements in $\mathfrak{R}$ have a size range from $\lambda+1$ to $3\lambda$.) For a successful binary search, however, it is sufficient to know that there is a $\lambda$ that satisfies the desired properties for step 3. Since we only enter this step if the extreme cases ($p$-BCD and 1-BCD) in step 1 and 2 hold, there has to exist at least one such $\lambda$ value in-between. By Lemma 3.10, a $\lambda$-BCD $(C_{2},H_{2},\mathfrak{R}_{2},f_{2})$ with $|H_{2}|+|\mathfrak{R}_{2}|>p$ implies $p_{\ell}>\lambda$. Since $p_{\ell}$ is an integer, it follows that $p_{\ell}\geq\lambda+1=c$.
For the size of the resulting instance, first note that a trivial no-instance has constant size. Otherwise, after step 5 the only component of size at most $c$ in $G$ is one copy of $K_{c}$. For the larger components, step 4 reduces further, if the $(p,c)$-CD is non-empty. By Lemma 3.13, a non-empty $(p,c)$-CD can be found if $|C_{1}|+|H_{2}|+|V(\mathfrak{R}_{2})|\geq p\max\{3c,p+c\}$. Thus exhaustive application of AlgVI terminates with a $c$-BCD $(C_{1},H_{1},\mathfrak{R}_{1},f_{1})$ that satisfies $|C_{1}|+|H_{2}|+|V(\mathfrak{R}_{2})|\leq p\max\{3c,p+c\}$ with a $c\leq p_{\ell}$. The union $C_{1}\cup H_{2}\cup V(\mathfrak{R}_{2})$ contains all vertices of the components in $G$ that have size larger than $c$, while after step 5 only one component of size at most $c$ remains. This gives a total of at most $p\max\{3c,p+c\}+c$ vertices in $G$. With $c\leq p_{\ell}<p$, if follows that the reduced graph contains at most $p(3(p-1))+p-1<3p^{2}$ vertices.
Finally, we prove the specified running time of Theorem 3, which completes the proof of Theorem 3. Note that for a yes-instance we have to call the algorithm AlgVI at most $|V|$ times to guarantee the desired kernel.
@@
Proof 3.18.
The calculation of a $\lambda$-BCD for $\lambda\in[p]$ runs in time $\mathcal{O}(|V|^{3}|E|)$ by Theorem 2, since a ($\lambda+1$)-packing has size less than $|V|$. The extraction of a $(p,\lambda)$-CD from it runs in time $\mathcal{O}(|V||E|)$ by Theorem 1. Each step consists of the calculation of these constructions with an additional time expenditure of $\mathcal{O}(\log(p))$ in step 3. As a result, the algorithm AlgVI runs in time $\mathcal{O}(\log(p)|V|^{3}|E|)$.
@@ 3.2 Weighted Vertex Integrity and Component Order Connectivity
In this section, we shift our focus to the weighted variants, namely wVI and wCOC. While utilizing the packing size of the associated BCD offers a starting point for deriving a lower bound for VI, it proves insufficient for improvements in the weighted setting. Therefore, we integrate the weight of the separator $H$ within a BCD $(C,H,\mathfrak{R},f)$ into our analysis. Additionally, we incorporate two distinct DBE’s into a BCD in case a reduction is not achievable within the respective setting. This approach enables us to establish a tighter lower bound for $p_{\ell}$, estimate the remaining instance size more accurately to obtain the desired kernelization results (cf. Theorems 4 and 5), or identify a no-instance.
We begin by introducing the two DBE structures mentioned in conjunction with a BCD, which serve as preconditions for the upcoming lemmas. Let $G=(V,E,w)$ be a vertex weighted graph, $a,s,\lambda\in\mathds{N}$ and $(C,H,\mathfrak{R},f)$ a $\lambda$-BCD with $\lambda\in[a]$ and $|H|+|\mathfrak{R}|\leq s$, where $\max_{v\in V}w(v)\leq\max(a,s)=:\mu$. Let $D^{Y}=\{d^{Y}_{h}\}_{h\in H}$ and $D^{Z}=\{d^{Z}_{h}\}_{h\in H}$ be demands, where $d^{Y}_{h}=a-2+\lambda\cdot(w(h)+1)$ and $d^{Z}_{h}=w(h)-1+(\sqrt{s}+1)\lambda$ for each $h\in H$. Let $(Y_{1},Y_{2},\lambda,f_{Y},D^{Y})$ and $(Z_{1},Z_{2},\lambda,f_{Z},D^{Z})$ be DBE’s in $G[C\cup H]$ with $Y_{1},Y_{2}$ as well as $Z_{1},Z_{2}$ are partitions of $H$ and $f_{Y}^{-1}(Y_{1}),f_{Y}^{-1}(Y_{2})$ as well as $f_{Z}^{-1}(Z_{1}),f_{Z}^{-1}(Z_{2})$ are partitions of $\mathbb{CC}(G[C])$.
The concept of two DBE’s embedded in a BCD can be understood as follows: With an appropriate choice of parameters $a$, $s$, and $\lambda$, the substructures $Y_{1}$ and $f_{Y}^{-1}(Y_{1})$ within the input graph of a wVI or wCOC instance form a reducible structure. Note that if $Y_{1}\neq\varnothing$ (or $Z_{1}\neq\varnothing$), these sets act as separators, isolating $f_{Y}^{-1}(Y_{1})$ (or $f_{Z}^{-1}(Z_{1})$) from the rest of the graph, as $H$ separates $C$ from the rest. However, if $Y_{1}=\varnothing$, attention shifts to $Z_{1}$. If the weight $w(Z_{1})$ exceeds a certain threshold, we can analyze the heaviest remaining component after removing a separator with bounded weight. If both $Y_{1}=\varnothing$ and $w(Z_{1})$ is bounded, we are able to bound the entire graph, or more precisely, all parts of the underlying $\lambda$-BCD $(C,H,\mathfrak{R},f)$.
Next, we present two crucial lemmas that serve as the foundation for designing the algorithms needed to prove Theorems 4 and 5. The first lemma aims to estimate the instance size.
@@
Let $w(Z_{1})\leq 2\mu^{1.5}$. If $Y_{1}=\varnothing$, then $w(\mathfrak{R})+w(C\cup H)<3\mu(s+\sqrt{\mu}\lambda)$.
Proof 3.20.
Let $C_{1}=V(f_{Z}^{-1}(Z_{1}))$ and $C_{2}=V(f_{Z}^{-1}(Z_{2}))$. We will use the demands $D^{Y}$ and $D^{Z}$ in the corresponding demanded balanced expansion to upper bound $w(Z_{1}\cup C_{1})$ and $w(Z_{2}\cup C_{2})$, respectively. Note that $C=C_{1}\cup C_{2}$ and $H=Z_{1}\cup Z_{2}$. Moreover, we use the properties of a $\lambda$-BCD to bound the weight of $R=\bigcup_{V^{\prime}\in\mathfrak{R}}V^{\prime}$ by $3|\mathfrak{R}|\lambda$, since $w(V^{\prime})\leq 3\lambda$ for every $V^{\prime}\in\mathfrak{R}$.
First, observe that $f_{Z}^{-1}(Z_{1})\subseteq f_{Y}^{-1}(Z_{1})$ since the components $f_{Z}^{-1}(Z_{1})$ are separated by $Z_{1}$ and must be assigned to $Z_{1}$ in $f_{Y}$. Note that $Y_{2}=H$ since $Y_{1}=\varnothing$, which implies that $Z_{1}\subseteq Y_{2}$. These facts allow us to bound $w(Z_{1}\cup C_{1})$ by the demands $D_{Y}$. As a result, we obtain
$w(Z_{1}\cup C_{1})$ $=\sum_{h\in Z_{1}}\left(w(h)+w(f_{Z}^{-1}(h))\right)$
$\leq\sum_{h\in Z_{1}}\left(w(h)+w(f_{Y}^{-1}(h))\right)$
$\leq\sum_{h\in Z_{1}}\left(d^{Y}_{h}+\lambda-1\right)$
$\leq\sum_{h\in Z_{1}}\left(a-2+\lambda\cdot(w(h)+1)+\lambda-1\right)$
$<\sum_{h\in Z_{1}}\left(a+2\lambda+w(h)\cdot\lambda\right).$
Furthermore, using the demands $D_{Z}$ we can bound $w(Z_{2}\cup C_{2})$, since for each $h\in Z_{2}$ we have
$w(h)+w(f_{Z}^{-1}(h))\leq d_{h}^{Z}+\lambda-1\leq w(h)-1+(\sqrt{s}+1)\lambda+\lambda-1<w(h)+2\lambda+\sqrt{s}\lambda.$
Now a straight forward calculation provides the proof:
$w(\mathfrak{R})+w(C\cup H)$ $\leq 3|\mathfrak{R}|\lambda+w(C_{1}\cup Z_{1})+w(C_{2}\cup Z_{2})$
$<3|\mathfrak{R}|\lambda+\sum_{h\in Z_{1}}(a+2\lambda+w(h)\cdot\lambda)+\sum_{h\in Z_{2}}(w(h)+2\lambda+\sqrt{s}\lambda)$
$=3|\mathfrak{R}|\lambda+|Z_{1}|(a+2\lambda)+|Z_{2}|(\sqrt{s}\lambda+2\lambda)+\lambda\cdot\sum_{h\in Z_{1}}w(h)+\sum_{h\in Z_{2}}w(h)$
$\leq 3|\mathfrak{R}|\lambda+(|Z_{1}|+|Z_{2}|)2\lambda+|Z_{1}|a+|Z_{2}|\sqrt{s}\lambda+w(Z_{1})\cdot\lambda+w(Z_{2})$
$<3|\mathfrak{R}|a+2|H|a+|Z_{1}|a+\mu^{1.5}\lambda+2\mu^{1.5}\lambda+|Z_{2}|\mu$
$\leq 3|\mathfrak{R}|\mu+2|H|\mu+(|Z_{1}|+|Z_{2}|)\mu+3\mu^{1.5}\lambda$
$\leq 3\mu(|\mathfrak{R}|+|H|)+3\mu^{1.5}\lambda$
$\leq 3\mu(s+\sqrt{\mu}\lambda)$
With respect to the weight of $Z_{1}$, the second lemma is relevant to Lemma 3.19 and helps to establish a more precise lower bound for $p_{\ell}$, or to identify instances where no feasible solutions exist.